# The effects of fibre spatial distribution and relative orientation on the percolation and mechanics of stochastic fibre networks: A model of peptide hydrogels

### v2

@@ Abstract
Researchers have modified various aspects of fibre networks to investigate their material properties. Parvez and Picu [11] investigated network connectivity – the mean number of fibre segments emerging from each node in the network (a segment is the portion of a fibre between two connected nodes). They showed that the elastic modulus in the non-affine regime increases, but the exponent $f_{h}$ decreases, with increasing connectivity. The global orientation preference of fibres has also attracted attention [12, 13]. Tarasevich and Eserkepov [12] modified fibre orientation with respect to a global coordinate system to create anisotropic networks and showed that they percolate at higher densities.
@@
In this paper, we model self-assembling peptide hydrogels (SAPHs), a family of peptides extensively studied by our group for their self-assembly, gelation, and potential biomedical uses like tissue engineering and drug delivery [18, 19, 20]. In SAPHs, fibres formed of short peptide sequences create a network through self-association [21]. The fibres are sufficiently thick and rigid enough to ignore thermal fluctuations. Variations in amino acid sequences or environmental conditions can significantly alter fibre interactions, affecting the network’s structure and mechanical properties, with elastic exponents ranging from 2 to 4.6 [10, 22]. Despite their importance, however, the effects of local fibre interactions on the structure and mechanical behaviour of the network are not fully understood. To understand how the structural changes caused by fibre interactions lead to mechanical changes, we modified the Mikado network generation process (we did not directly model fibre formation, branching and interaction). Our results apply equally to any materials that have similar structures.
@@
where $P_{r}$, $n_{e}$, and $\alpha$ are, respectively: the reduction in the probability of the new fibre existing due to the presence of a single fibre centre previously placed within the circle, the number of fibre centres that already exist within that circle, and a constant which scales $r_{e}$ relative to $\sqrt{1/\rho}$, the distance between nodes in perfect square lattice. In our simulations we chose $\alpha=3/4$. The orientation is chosen, based on the orientations of the existing fibres within a circle of radius $r_{\theta}$ (we chose $r_{\theta}=l/2)$, from a Gaussian mixture distribution [23]:
@@
where $\Theta$, $n_{\theta}$, $\sigma$, and $\theta_{fi}$, are the preferred orientation relative to the existing fibres, the number of fibres within the circle, the standard deviation of the distribution (we used $\sigma=1/\sqrt{12}$), and the angle the $i^{\text{th}}$ fibre makes with the $x$-axis, respectively.
@@
Figure 1: Samples of the studied networks with $\rho=30$, $l=1$, and $L=5$.
@@
The second aspect that we investigated was how fibres are dispersed within the domain of study. To do so, we created networks of various densities with $L=10$, divided them into 400 smaller boxes, counted the number of fibre centres in each box, then calculated the mean and variance of this quantity. It is well known that, in Mikado networks, the number of fibre centres in each box follows a Poisson distribution [28] (Figure 2(a)). An increase in the value of $P_{r}$ results in a more homogeneous (lower variance) network.
@@
The connectivity of a network, $Z$, plays an important role in the network’s behaviour [11]. As depicted in Figure 2(b), for more homogeneous networks, the connectivity is slightly lower. The reason for this is fibre clumping (see Figure 1) [28]. As a result, fibres in the vicinity of these clumps have a relatively high chance of forming connections, and thus will have a higher connectivity. Moreover, for the networks whose fibres have a preference for being parallel ($\Theta=0$), the connectivity is considerably smaller. However, the difference between all network classes vanishes as the density increases.
@@
Our investigation into the networks’ mechanical properties is divided into the two cases of near percolation and high density [2, 7]. For low densities, the finite size scaling method [8, 7] is used, and at high densities, a representative volume element (RVE) is considered [9]. In this study, the segments of the network are considered to be beams with second moment of area, $I$ [29], with square cross-sections. The nodes are considered to be stiff cross-links (relative rotation is prohibited between beams connected at a node). The Timoshenko beam model [29] is used to model the beams and the direct stiffness method [29] is used to model infinitesimal deformation of the network.
@@
To calculate the elastic modulus, a displacement in the $y$-direction was applied to each node on the top boundary, while the displacement in the $x$-direction and rotation were fixed. On the bottom boundary, zero displacement and rotation was imposed, and on the lateral boundaries, zero force was imposed. The number of realisations varied with system size from 1,600 for the largest to 102,600 for the smallest.
@@
The results are presented in Table 1. For the Mikado networks, the value of the elastic exponent was calculated as $f=3.73\pm 0.02$. This is different from a previously reported value, $f=3.2\pm 0.3$ [7, 2], due to our use of a more accurate value for the percolation threshold. Our results show that the various network types considered here have different elastic exponents. This suggests that materials whose fibres interact during formation, thus modifying the structure of the network, belong to different universality classes.
@@
It has been shown that, at high density, the behaviour of fibre networks can be divided into two regimes: affine and non-affine [9, 2, 7]. In the non-affine regime, the bending stiffness of the fibre segments is much smaller than their stretching stiffness. In the affine regime, the network elastic moduli vary linearly with density and in the non-affine regime, they vary proportionally to $(\rho l)^{f_{h}}(l_{b}/l)^{2}$, where $l_{b}=\sqrt{I/A}$, and $A$ is the fibre cross-sectional area [9].
To determine the shear modulus, $G$, of the simulated networks, a periodic boundary condition with a far-field simple shear deformation was applied in a manner that is explained in depth by Nguyen et al. [31]. In order to find $f_{h}$ for different network types, we simulated networks with a fibre thickness of 0.0001 and densities ranging from 25 to 500; the results are presented in Figure 3(a). For the Mikado network, $f_{h}$ was found to be 7.9, as reported in the literature [9, 32]. The other values are presented in Table 1. Increasing the spatial dispersion of fibres slightly decreases the value of $f_{h}$. Furthermore, noticeable changes can be observed for networks with preferred parallel orientations. It is likely that all of the geometrical features discussed in Figure 2 impact the network mechanical moduli, but disentangling every specific relationship between structure and mechanics is not straightforward. However, as can be seen from Figures 2b-d, for DOM type 3 and 4 networks, the connectivity is lower, and the gradients of the mean segment length and the mean segment density are steeper; which, in the bending-dominated regime, causes the value of $f_{h}$ to be higher.
@@
The value of $G$ varies significantly between the six network types. It can vary up to 50 times in the bending-dominated regime. For DOM type 3 and 4 networks, $G$ is smaller by at least one order of magnitude in the non-affine regime. This is due to their lower connectivity and segment density, and higher segment length. This clearly shows how influential network structure is on the mechanical behaviour of fibrous materials.
Figure 3: (a) Shear modulus as a function of density for high density networks ($E_{f}$ is the fibre Young’s modulus). (b) The dependence of Poisson’s ratio on density and $l_{b}$. For the main figure $l_{b}=2.9\mathrm{e}{-5}$. In the inset panel, $\rho=75$. (c) $PS_{n\%}$ represents the smallest possible proportion of a network’s segments that together store $n\%$ of the total energy in the network. $PS_{n\%}^{\rho=25}$ represents $PS_{n\%}$ with $\rho=25$. All the data is this figure is for Mikado networks in the non-affine regime. (d) Histogram of the orientations of fibres relative to the horizontal axis (which is perpendicular to the stretch direction) that carry $50\%$ of the energy in each Mikado network. The absolute value of each angle has been taken. This data is for $l_{b}=2.9\mathrm{e}{-6}$.
@@
We also investigated the Poisson’s ratio, $\nu$, of our six network types. To do so, we applied boundary conditions that impose a pure longitudinal deformation, and calculated the total forces in the directions parallel and perpendicular to the direction of stretch; Poisson’s ratio is then calculated using $\nu=F_{\perp}/F_{\parallel}$, where $F_{\perp}$ and $F_{\parallel}$ are the total perpendicular and parallel forces at the boundary interfaces (see Supplemental Material for a derivation). We display the effects of changing the density, $\rho$, and $l_{b}$, in Figure 3(b). There is a noticeable difference in the value of $\nu$ between the different networks, which is more pronounced when the relative orientations of the fibres change.
@@
The Poisson’s ratio in the non-affine regime is not constant and changes with increasing density. This is contrary to previous papers that reported $\nu$ to be constant [2, 32]. The effect on $\nu$ of increasing density diminishes into the affine regime (as far as we have made observations). The effect of $l_{b}$, however, is different; in the non-affine regime, increasing it has no effect, but as we transition into the affine regime, increasing it reduces the value of $\nu$ as far as we have made observations.
The reason for the effect of density on $\nu$, even when the network remains in the non-affine regime, lies in how the network absorbs energy at different densities. We have demonstrated this for Mikado networks by looking at two parameters: the percentage of segments that carry a given percentage of the total energy absorbed by the network, and the orientations of these fibres. The results are presented in Figures 3(c) and 3(d). The percentage of segments involved in absorbing a given portion of the total energy decreases with increasing density. The orientation distribution of these fibres also noticeably changes. In Figure 3(d), we can see that, at higher densities, the segments have a greater preference to be parallel to the stretch direction (in all cases the energy absorbed through stretching is negligible). The reason for the change in the energy distribution is unclear, but it can be inferred that this change is rooted in how the microstructure deforms and this, in turn, affects the Poisson’s ratio.
To compare the results of the model networks qualitatively with real SAPHs, we carried out shear rheometry tests on three $\beta$-sheet forming peptides: KFEFEFKFKF (F), KFEFEFKFKK (K), and KFEFEFKFKFK (FK) (K: lysine; F: phenylalanine; E: glutamic acid). These peptides are based on a design first developed by Zhang et al. [33], Zhang and Altman [34] (see the Supplemental Material for details). Figure 4a-c shows transmission electron microscopy (TEM) images obtained for 10$\times$ diluted hydrogels alongside the chemical structure of the SAPHs. These images confirm the formation in all three systems of semi-rigid peptide fibre networks with the thinnest fibre diameter ranging from 3 to 4 nm. Wychowaniec et al. [22] have shown that peptides with hydrophobic end-groups (F) have more hydrophobic edge interactions and are thought, therefore, to form more crosslinks. These networks are thus more similar to the perpendicular network model described above. On the other hand, peptides with hydrophilic (K) end-groups were shown to make fewer crosslinks due to edge electrostatic repulsion and, therefore, will be more similar to the parallel network model. In Figure 4d, the storage moduli, $G^{\prime}$, obtained for these three hydrogels are presented. For FK and K hydrogels, similar exponents (2.4 $\pm$ 0.2) were obtained. Both these systems have fibres with hydrophilic edges due to the presence of the terminal K. FK was found to have a higher $G^{\prime}$ compared to K which can be attributed to the presence of the additional phenylalanine group before the terminal K. This additional F increases the overall hydrophobicity of the peptides and the intermolecular forces (H-bond number) between peptides, resulting in more rigid fibres with higher moduli, without affecting the exponent. For the F hydrogels, significantly higher values of $G^{\prime}$ were observed (due to higher crosslinking levels) with a smaller exponent, as predicted by the models.

### v3

@@ 1 Introduction
Researchers have modified various aspects of fibre networks to investigate their material properties. Parvez and Picu [11] investigated network connectivity – the mean number of fibre segments emerging from each node in the network (a segment is the portion of a fibre between two connected nodes). They showed that the elastic modulus in the bending dominated regime increases, but the exponent $f_{h}$ decreases, with increasing connectivity (average number of connections to other nodes that each network node has). The global orientation preference of fibres has also attracted attention [12, 13]. Tarasevich and Eserkepov [12] modified fibre orientation with respect to a global coordinate system to create anisotropic networks and showed that they percolate at higher densities.
@@
In this paper, we model self-assembling peptide hydrogels (SAPHs), a family of peptides extensively studied by our group for their self-assembly, gelation, and potential biomedical uses such as tissue engineering and drug delivery [18, 19, 20]. In SAPHs, fibres formed of short peptide sequences create a network through self-association [21]. The fibres are sufficiently thick and rigid to ignore thermal fluctuations. Variations in amino acid sequences or environmental conditions can significantly alter fibre interactions, affecting the network’s structure and mechanical properties, with elastic exponents ranging from 2 to 4.6 [10, 22]. Despite their importance, however, the effects of local fibre interactions on the structure and mechanical behaviour of the network are not fully understood. To understand how the structural changes caused by fibre interactions lead to mechanical changes, we modified the Mikado network generation process (we did not directly model fibre formation, branching, interaction, filament entanglements, and thermal fluctuations) to vary the fibres’ spatial distribution and local orientations relative to each other. Our results can be applied to any materials that are discovered, or developed, to have similar features.
2 Model
@@
where $P_{r}$, $n_{e}$, and $\alpha$ are, respectively: the reduction in the probability of the new fibre existing due to the presence of a single fibre centre previously placed within the circle, the number of fibre centres that already exist within that circle, and a constant which scales $r_{e}$ relative to $\sqrt{1/\rho}$, the distance between nodes in perfect square lattice. In our simulations we chose $\alpha=4/3$, which maximises the homogeneity of the fibres’ spatial distibution (for smaller and larger values, the fibre density is less homogeneous). The orientation is chosen, based on the orientations of the existing fibres within a circle of radius $r_{\theta}$ (we chose $r_{\theta}=l/2$, since this is the maximal distance between two fibre centres such that they can theoretically be connected by a fibre), from a Gaussian mixture distribution [23]:
@@
where $\Theta$, $n_{\theta}$, $\sigma$, and $\theta_{fi}$, are the preferred orientation relative to the existing fibres, the number of fibres within the circle, the standard deviation of the distribution (we used $\sigma=1/\sqrt{12}$, a value small enough to ensure the angle of the new fibre is sufficiently affected by existing fibres), and the angle the $i^{\text{th}}$ fibre makes with the $x$-axis, respectively.
@@
Figure 1: Samples of the studied networks with $\rho=30$, $l=1$, and $L=5$. In the Mikado network, the red arrows indicate high density clusters.
3 Results
3.1 Geometrical features
@@
The second aspect that we investigated was how fibres are dispersed within the domain of study. To do so, we created networks of various densities with $L=10$, divided them into 400 smaller boxes, counted the number of fibre centres in each box, then calculated the mean and variance of this quantity. It is well known that, in Mikado networks, the number of fibre centres in each box follows a Poisson distribution [28] (Figure 2(a)). An increase in the value of $P_{r}$ results in a more homogeneous (lower variance) network (with fewer high density clusters, as indicated in the Mikado network in Figure 1).
@@
It has been shown that the connectivity of a network, $Z$, plays an important role in the network’s behaviour [11]. As depicted in Figure 2(b), for more homogeneous networks, the connectivity is slightly lower. The reason for this is fibre clustering (see Figure 1) [28]. As a result, fibres in the vicinity of these clusters have a relatively high chance of forming connections, and thus will have a higher connectivity. Moreover, for the networks whose fibres have a preference for being parallel ($\Theta=0$), the connectivity is considerably smaller. However, the difference between all network classes vanishes as the density increases.
@@
3.2 Mechanical behaviour
Our investigation into the networks’ mechanical properties is divided into the two cases of near percolation and high density [2, 7]. For low densities, the finite size scaling method [8, 7] is used, and at high densities, a representative volume element (RVE) is considered [9]. In this study, the segments of the network are considered to be beams with second moment of area, $I$ [29], with square cross-sections. The nodes are considered to be stiff cross-links (relative rotation is prohibited between beams connected at a node). The Timoshenko beam model [29] is used to model the beams and the direct stiffness method [29] is used to model infinitesimal deformation of the network. In all of our simulations, the length of each fibre is one, the thickness and density are varied as mentioned in each section, and the reported moduli are normalized on the fibres’ Young’s modulus.
@@ 3.2 Mechanical behaviour
To calculate the elastic modulus, a displacement in the $y$-direction was applied to each node on the top boundary, while the displacement in the $x$-direction and rotation were fixed. On the bottom boundary, zero displacement and rotation was imposed, and on the lateral boundaries, zero force was imposed. The number of realisations varied with system size from 1,600 for the largest to 102,600 for the smallest (see the Supplemental Material for further details).
@@
The results are presented in Table 1. For the Mikado networks, the value of the elastic exponent was calculated as $f=3.73\pm 0.02$. This is different from a previously reported value, $f=3.2\pm 0.3$ [7, 2], due to our use of a more accurate value for the percolation threshold. Our results show that the various network types considered here have different elastic exponents. This suggests that materials whose fibres interact during formation, thus modifying the structure of the network, have mechanical behaviour that scales differently with density to materials whose fibres do not interact. The value of the elastic exponent near percolation specifies how the material stiffens. The value of $f$ is especially important for materials whose fibre density is close to the percolation density, such as particulate gels. It has recently been discussed by Bantawa et al. [31] that one possible reason for the difference in the observed exponent in colloidal gels is the fractal nature of the edges in the structure. They introduced an equation that relates the fractal dimensionality of the strains to the observed exponent: $f_{obs}=f/\nu_{c}(d-d_{f})$, where $f_{obs}$, $\nu_{c}$, $d$, and $d_{f}$ are the observed exponent, correlation exponent, dimensionality, and fractal dimensionality of the gel chains, respectively. However, they considered the exponent $f$ to be constant for all of their systems. We have shown that $f$ changes when the network structure changes in the ways described above.
@@
It has been shown that, at high density, the behaviour of fibre networks can be divided into two regimes: bending dominated and stretching dominated [9, 2, 7] (we have defined the border of the bending dominated regime as being the density threshold below which more than $99\%$ of the total energy in the network is stored in bending). In the bending dominated regime, the bending stiffness of the fibre segments is much smaller than their stretching stiffness. In the stretching dominated regime, the network elastic moduli vary linearly with density and in the bending dominated regime, they vary proportionally to $(\rho l)^{f_{h}}(l_{b}/l)^{2}$, where $l_{b}=\sqrt{I/A}$, and $A$ is the fibre cross-sectional area [9].
To determine the shear modulus, $G$, of the simulated networks, a periodic boundary condition with a far-field simple shear deformation was applied in a manner that is explained in depth by Nguyen et al. [32]. In order to study $G$ for different network types, we simulated networks with fibre thicknesses ranging from $10^{-5}$ to $10^{-3}$ and densities ranging from 25 to 500; the results are presented in Figure 3. In Figure 3(a), we used several different values of $l_{b}$ corresponding to fibre thicknesses in the stated range (by plotting $G/\rho E_{f}A$, where $E_{f}$ is the fibre Young’s modulus, as a function of $(\rho l)^{f_{h}-1}(l_{b}/l)^{2}$, the data for each network type collapse onto a single curve called a master curve by some authors), whereas in Figure 3(b), we used a fixed fibre thickness of $10^{-4}$.
Figure 3: (a) Master curves for the (non-dimensionalised) shear modulus. The values of $f_{h}$ are reported in Table 1. The arrows indicate the boundary of the bending dominated regime. (b) (Non-dimensionalised) shear modulus as a function of density for high density networks.
@@
The value of $G$ varies significantly between the six network types. It can vary up to 50 times in the bending-dominated regime (peptide hydrogels exist mainly in this regime). For DOM type 3 and 4 networks, $G$ is smaller by at least one order of magnitude in the bending dominated regime. This is due to their lower segment density, and higher segment length, and, for the same reason, the transition away from the bending dominated regime occurs later. This clearly shows how influential network structure is on the mechanical behaviour of fibrous materials.
For the Mikado network, $f_{h}$ was found to be 7.9, as reported in the literature [9, 33]. The other values are presented in Table 1. Increasing the spatial dispersion of fibres slightly decreases the value of $f_{h}$. Furthermore, noticeable changes can be observed for networks with preferred parallel orientations. All of the geometrical features discussed in Figure 2 impact the network mechanical moduli to some degree, but disentangling every specific relationship between structure and mechanics is not straightforward. It has been shown that increasing the connectivity affects the elastic moduli [11], but the changes in connectivity that we observe here (Figure 2(b)) are very small compared to those studied by Parvez and Picu [11]. For DOM type 3 and 4 networks, the slope of the mean segment length and the mean segment density are steeper; which, in the bending-dominated regime, causes the value of $f_{h}$ to be higher.
@@
We also investigated the Poisson’s ratio, $\nu$, of our six network types. To do so, we applied boundary conditions that impose a pure longitudinal deformation, and calculated the total forces in the directions parallel and perpendicular to the direction of stretch; Poisson’s ratio is then calculated using $\nu=F_{\perp}/F_{\parallel}$, where $F_{\perp}$ and $F_{\parallel}$ are the total perpendicular and parallel forces at the boundary interfaces (see Supplemental Material for a derivation). We display the effects of changing the density, $\rho$, and $l_{b}$, in Figure 4(a). There is a noticeable difference in the value of $\nu$ between the different networks, which is more pronounced when the relative orientations of the fibres change.
@@
The Poisson’s ratio in the bending dominated regime is not constant and changes with increasing density. This is contrary to previous papers that reported $\nu$ to be constant [2, 33]. The effect on $\nu$ of increasing density diminishes into the stretching dominated regime (as far as we have made observations). The effect of $l_{b}$, however, is different; in the bending dominated regime, increasing it has no effect, but as we transition into the stretching dominated regime, increasing it reduces the value of $\nu$ as far as we have made observations.
Figure 4: (a) The dependence of Poisson’s ratio on density and $l_{b}$. For the main panel, $l_{b}=2.9\mathrm{e}{-5}$. In the inset panel, $\rho=75$.(b) A snapshot of the Mikado network with density 25 and $L=15$. The red segments represent the smaller portion of the network that carries 50% of the total energy of the system and the blue segments represent the largest possible portion of the network that only carries 5% of the total energy. (c) A snapshot of the network with $\rho=60$ and $L=5$. (d) The average length of the segments that carry 50% of the total energy, $\overline{l_{c}^{50\%}}$, divided by average segment length $\overline{l_{c}}$. (e) Histogram of node rotation, $\omega$, at two different densities at a strain of $1/1000$. (f) The average absolute value of rotation versus density.
As the density increases, how the network deforms changes even though we are still in the bending dominated regime – this can be seen by examining how energy is stored in the network and how the nodes rotate (Figures 4(b) and 4(c)). At smaller densities, the low density areas between clusters absorb energy (undergo deformation) and the clusters, where the density of segments is high and the average lengths are shorter, do not deform significantly. This means that these dense clusters move (rotate/translate) almost rigidly (with a low average contribution to absorbing energy). It can be observed, by looking at Figure 4(d), that, at low density, the average length of the segments that absorb most of the energy is higher than the average segment size of the network. As the density increases, this value decreases, and more fibres within the clusters contribute to absorbing energy. Furthermore, as the density increases, more segments deform (due to node rotation), as can be seen from Figure 4(e) (note that the rotation of all nodes is zero for an affine deformation). The average absolute rotation is also plotted in Figure 4(f), which shows that the same trend is followed by all network classes. Since the deformation of the microstructure changes as we increase the density, we may expect a change in the Poisson’s ratio (which is a direct result of microstructural movement). The low density deformation is greatly affected by the microstructure and the Poisson’s ratios are noticeably different for different network classes as a result. At higher densities, however, the Poisson’s ratio values converge since, at higher densities, the interiors of the clusters also deform and, therefore, the effects of microstructure decrease gradually.
3.3 Experiments
To compare the results of the model networks qualitatively with real SAPHs, we carried out shear rheometry tests on three $\beta$-sheet forming peptides: KFEFEFKFKF (F), KFEFEFKFKK (K), and KFEFEFKFKFK (FK) (K: lysine; F: phenylalanine; E: glutamic acid). These peptides are based on a design first developed by Zhang et al. [34], Zhang and Altman [35] (see the Supplemental Material for details). Figure 5a-c shows transmission electron microscopy (TEM) images obtained for 10$\times$ diluted hydrogels alongside the chemical structure of the SAPHs. These images confirm the formation in all three systems of semi-rigid peptide fibre networks with the thinnest fibre diameter ranging from 3 to 4 nm. Considering the processing required, TEM images cannot give us completely accurate information about the in situ structure of the networks directly. To obtain these images, we dilute the original hydrogel 10 times and then dry it on a grid. Due to this processing, some structural information is lost. However, we can conclude the following important points from the TEM images: 1) the fibres do not aggregate – they are the same size in each network, which rules out mechanical effects that might arise from aggregation; 2) almost all of the peptides form fibres and they connect to each other (if this was not the case, a direct relation between density and peptide concentration would not exist); 3) the fibres form in approximately straight lines.
@@ 3.3 Experiments
Wychowaniec et al. [22] have shown that peptides with hydrophobic end-groups (F) have more hydrophobic edge interactions and are thought, therefore, to form more crosslinks. These networks are thus more similar to the perpendicular network model described above. On the other hand, peptides with hydrophilic (K) end-groups were shown to make fewer crosslinks due to edge electrostatic repulsion and, therefore, will be more similar to the parallel network model. In Figure 5d, the storage moduli, $G^{\prime}$, obtained for these three hydrogels are presented. For FK and K hydrogels, similar exponents (2.4 $\pm$ 0.2) were obtained. Both these systems have fibres with hydrophilic edges due to the presence of the terminal K. FK was found to have a higher $G^{\prime}$ compared to K which can be attributed to the presence of the additional phenylalanine group before the terminal K. This additional F increases the overall hydrophobicity of the peptides and the intermolecular forces (H-bond number) between peptides, resulting in more rigid fibres with higher moduli, without affecting the exponent. For the F hydrogels, significantly higher values of $G^{\prime}$ were observed (due to higher crosslinking levels) with a smaller exponent, as predicted by the models.
It has been observed that models with different dimensionality have qualitatively similar behaviour [36] but different exponents (the exponents of Mikado networks in 2D and 3D are 3 and 8, respectively [1]). Despite the difference in the actual exponent observed in our model and in our experiments, the trends in the behaviour of the exponents are similar, and could arise from a similar phenomenon (the difference in the values originates from the difference in dimensionality, which is expected). For example, the attraction and repulsion forces between fibres would affect the structure of the network, and locally rearrange it, regardless of whether the system is 2D or 3D.
4 Conclusion